This question is about LCM and GCD:
A rectangular field with dimensions 75 ft. by 625 ft. is to be divided into same-size square plots. If the sides of the squares need to be whole numbers of feet long, what are the largest squares possible and how many such squares will fit in the field? Please help :)

Answer by jsmallt9(3759)

The area of this rectangular field is 75*625. Since 625 = 25*25, 75*625 = 75*25*25. From this we can see that 75 25 by 25 squares will fit.